In 10 years, a man will be 3 times as old as his son is now. If the son's present age is x years, find the father's present age.
Answer by JulietG(1812)

You can't solve this fully without an actual value, but here's the way to do it.
Let m be dad and x be son
m+10 = 3x
.
Let's say the son is 12.
m+10 = 3*12
m = 36-10 or 26 (highly improbable!)
.
The son is more likely older; let's say 20
m+10 = 60
m = 60 - 10 = 50
That makes more sense.